A body weighs Win a liquid of density p1 and Win a liquid of densith p2. What is the weight of body in a liquid of density p3. 


p = rho


Grade:

1 Answers

vikas askiitian expert
13 years ago

let actual weiight of body is W & volume is V then

 W - p1V.g = W1       ........1

 W - p2Vg = W2     ........2

solving  1 & 2

 W =  (W1p2-W2p1)/p2-p1        &         V = (W1-W2)/g(p2-p1)

 

now , let its weight in liquid P3 is W3 then

 W - p3V.g = W3

now putting V & W ,

 we get

 W3 = [W1p2-W2p1-p3(W1-W2)]/(p2-p1)

approve my ans if u like it